# publications.bib

### v3

@@ 3. Mantovan’s Formula and the Hodge-Tate Period Morphism
This can be checked using the moduli interpretation, as in the proof of [CS2, Theorem 4.4.1]. In particular, given a $\mathop{\rm Spa}(C^{\prime},\mathcal{O}_{C^{\prime}})$ point of $\mathfrak{Ig}^{b,\mathrm{tor}}_{C}$ specializing to a boundary component indexed by an Igusa cusp label $\tilde{Z}$ (See §2.2.2 for the definition of Igusa cusp label), then from the discussion in §2.2.3 this corresponds to the datum of $(\mathcal{B},\mathcal{G},\mathcal{G}^{\vee},\rho,y)$, where
  1. (1)
    $\mathcal{B}$ is an abelian scheme over $\mathcal{O}_{C^{\prime}}$ with polarization, and $\mathcal{O}_{B}$-structure,
  2. (2)
    $\mathcal{G}$ is a Raynaud extension
    of $\mathcal{B}$ by a torus $T$ with cocharacter group $X$,
  3. (3)
    $\mathcal{G}^{\vee}$ is a Raynaud extension
    of $\mathcal{B}^{\vee}$ by a torus $T^{\vee}$ with cocharacter group $Y$,
  4. (4)
    $\rho$ is an $\mathcal{O}_{B}$-linear isomorphism $\rho:\mathcal{G}[p^{\infty}]\simeq Z_{b,-1}$, extending the isomorphism $T[p^{\infty}]\simeq Z_{b,-2}$,
  5. (5)
    $y\in\mathcal{P}_{\Sigma_{\Phi}}(\mathcal{O}_{C^{\prime}})$, where $\mathcal{P}_{\Sigma_{\Phi}}$ is the toroidal compactification (determined by an admissible rpcd $\Sigma_{\Phi}$, as in §2.1.2) of a torsor under a torus $E_{\Phi}/\mathcal{O}_{C^{\prime}}$, whose character group is given by $S_{\Phi}$, and $y$ is a point whose special fiber lies in the boundary.
The natural map $\mathfrak{Ig}^{b,\mathrm{tor}}(C^{\prime},\mathcal{O}_{C^{\prime}})\hookrightarrow\mathcal{S}_{K^{p},C}^{\mathrm{tor}}(C^{\prime},\mathcal{O}_{C^{\prime}})$ is determined by forgetting the trivialization $\rho:Z_{b,-1}\simeq\mathcal{G}[p^{\infty}]$. The image under this map of a point lying in the boundary $\partial\mathfrak{Ig}^{b,\mathrm{tor}}_{C}$ is equivalent to insisting that $y$ is a point whose special fiber lies in the boundary of $\mathcal{P}_{\Sigma_{\Phi}}$ for some Igusa cusp label $\tilde{Z}$, which is not the trivial one. Let $Z=(Z,\Phi)$ be the cusp label which $\tilde{Z}$ lives over. This is then equivalent to the condition guaranteeing that the image lies in the component of $\mathcal{S}_{K^{p},C}^{\mathrm{tor}}$ indexed by $Z=(Z,\Phi)$, where $Z$ is not the trivial cusp label, and thus lies in the boundary $\partial\mathcal{S}^{\mathrm{tor}}_{K^{p},C}$. The claim follows.
@@
For a geometric point $x:\mathop{\rm Spa}(C,C^{+})\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{b}$, we have an identification:
@@
We have an identification $\pi_{\mathrm{HT}}^{-1}(x)\simeq\overline{\mathfrak{Ig}}^{b,*}_{C}\setminus\overline{\partial\mathfrak{Ig}}^{b,*}_{C}$ by the previous Corollary, so, by proper base-change, we are tasked with computing the compactly supported cohomology of this space. We note, by Theorem 3.3, the adic spaces $\mathfrak{Ig}^{b,*}_{C}$ are affinoid perfectoid. It follows that the canonical compactification $\overline{\mathfrak{Ig}}^{b,*}_{C}\simeq(\pi_{\mathrm{HT}}^{*})^{-1}(x)$ is also affinoid perfectoid, by [Ecod, Proposition 18.7 (iv)]. In particular, it is quasi-compact and partially proper, so in particular proper. It therefore follows by excision [^1] One easily checks that the excision sequence is exact on points, and this is sufficient by [Ecod, Proposition 14.3]. that we have a distinguished triangle
@@
Applying Theorem 3.2 again, we know that $k:\mathfrak{Ig}^{b,*}\hookrightarrow\overline{\mathfrak{Ig}}^{b,*}_{C}$ is a qcqs open immersion of perfectoid spaces inducing an isomorphism on rank $1$ points, and it follows that the same is true for the induced map on the Zariski closed subspaces $\phantom{}{}^{\partial}k:\partial\mathfrak{Ig}^{b,*}\hookrightarrow\overline{\partial\mathfrak{Ig}}^{b,*}_{C}$. Therefore, we can apply [CS1, Lemma 4.4.2], this tells us that the natural maps
@@
are isomorphisms, giving identifications $R\Gamma(\overline{\mathfrak{Ig}}_{C}^{b,*},\overline{\mathbb{F}}_{\ell})\simeq R\Gamma(\mathfrak{Ig}^{b,*}_{C},\overline{\mathbb{F}}_{\ell})$ and $R\Gamma(\overline{\partial\mathfrak{Ig}}_{C}^{b,*},\overline{\mathbb{F}}_{\ell})\simeq R\Gamma(\partial\mathfrak{Ig}^{b,*}_{C},\overline{\mathbb{F}}_{\ell})$. Now, by [CS1, Lemma 4.4.3], we have further identifications of $R\Gamma(\partial\mathfrak{Ig}^{b,*}_{C},\overline{\mathbb{F}}_{\ell})$ and $R\Gamma(\mathfrak{Ig}^{b,*}_{C},\overline{\mathbb{F}}_{\ell})$ with the cohomology of the perfect schemes $\partial\mathrm{Ig}^{b,*}$ and $\mathrm{Ig}^{b,*}$, respectively. Substituting this into the triangle (3), we get a distinguished triangle
@@
By applying quasi-compact base-change [Ecod, Proposition 17.6] and then using that the inclusion $\partial\mathfrak{Ig}_{C}^{b,*}\subset\mathfrak{Ig}_{C}^{b,*}$ is induced from taking the rigid generic fiber over $C$ of Witt vectors applied to $\partial\mathrm{Ig}^{b,*}\subset\mathrm{Ig}^{b,*}$, we identify the last map with the natural restriction map on the cohomology. However, this identifies the first term with precisely the partially compactly supported cohomology, as desired. ∎
@@
If $d_{b}:=\langle 2\rho_{G},\nu_{b}\rangle=\mathrm{dim}(\mathrm{Ig}^{b,*})=\mathrm{dim}(\mathrm{Ig}^{b})$ then the cohomology of the complex
@@
We saw in Theorem 3.3 that $\mathrm{Ig}^{b,*}$ is an affine scheme. So we would like to apply Artin vanishing; however, $\mathrm{Ig}^{b,*}$ is also a perfect scheme so in particular not of finite type. To remedy this, consider the pro-étale cover
$\mathrm{Ig}^{b}\rightarrow\mathcal{C}_{\mathbb{X}_{b}}^{\mathrm{perf}},$
with Galois group $\mathrm{Aut}(\mathbb{X})(\overline{\mathbb{F}}_{p})$ over the perfection of the central leaf attached to $\mathbb{X}_{b}$. This is obtained as the perfection of the limit of the finite étale covers
$\mathrm{Ig}^{b}_{m}\rightarrow\mathcal{C}_{\mathbb{X}_{b}},$
described in 2.2.3, as shown in [CS1, Proposition 4.3.8]. These spaces are of finite type over $\overline{\mathbb{F}}_{p}$. We now define $\mathrm{Ig}^{b,*}_{m}$ to be the normalization of $\mathcal{C}_{\mathbb{X}}^{*}$ in $\mathrm{Ig}^{b}_{m}$ of the finite étale cover $\mathrm{Ig}^{b}_{m}\rightarrow\mathcal{C}_{\mathbb{X}_{b}}$. By [San, Theorem 3.33], $\mathcal{C}_{\mathbb{X}}^{*}$ is affine; therefore, it follows that $\mathrm{Ig}^{b,*}_{m}$ is a normal and affine scheme which will be of finite type, since $\mathrm{Ig}^{b}_{m}$ is. It follows, by definition of $\mathrm{Ig}^{b,*}$, that it is the perfection of $\mathop{\rm lim}\nolimits_{m\geq 1}\mathrm{Ig}_{m}^{b,*}$. Therefore, since passing to perfections doesn’t change the étale cohomology, we can conclude by combining Artin vanishing with an application of [stacks-project, Tag 09QY] to the system of sheaves $g_{b,m!}(\overline{\mathbb{F}}_{\ell})$, where $g_{b,m}:\mathrm{Ig}^{b}_{m}\rightarrow\mathrm{Ig}^{b,*}_{m}$ is the natural open inclusion at finite level. ∎
Now we would like to link this analysis with the semi-perversity of certain sheaves on $\mathrm{Bun}_{G}$. We consider the Hodge-Tate period morphism
$\pi_{\mathrm{HT}}:[\mathcal{S}_{K^{p}}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]$
quotiented out by $G(\mathbb{Q}_{p})$. We let $h^{\rightarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\underline{G(\mathbb{Q}_{p})}]\simeq\mathrm{Bun}_{G}^{1}$ be the structure map quotiented out by $G(\mathbb{Q}_{p})$. Note this is a proper map, since $\mathcal{F}\ell_{G,\mu^{-1}}$ is proper over $\mathop{\rm Spd}(C)$.
@@
(4) $R\Gamma_{c}(\mathcal{S}_{K^{p},C},\overline{\mathbb{F}}_{l})\simeq h^{\rightarrow}_{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$
of $G(\mathbb{Q}_{p})$-representations, and this computes the compactly supported torsion cohomology of the Shimura variety.
@@
$h^{\leftarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]\rightarrow\mathrm{Bun}_{G}$
@@
remembering the isomorphism class of the bundle $\mathcal{E}_{1}$ in the moduli interpretation of $\mathcal{F}\ell_{G,\mu^{-1}}$ as a diamond. This defines a cohomologically smooth map by [FS, Theorem IV.1.19], and the image identifies with the open subset $B(G,\mu)\subset B(G)$ under the identification $|\mathrm{Bun}_{G}|\simeq B(G)$ of topological spaces, where $|\mathrm{Bun}_{G}|$ denotes the underlying topological space of $\mathrm{Bun}_{G}$ and $B(G)$ has the topology given by its natural partial ordering [Vi]. For each $b$, we have a locally closed Harder-Narasimhan stratum $j_{b}:\mathrm{Bun}_{G}^{b}\hookrightarrow\mathrm{Bun}_{G}$, and we can define the locally closed subset $[\mathcal{F}\ell_{G,\mu^{-1}}^{b}/\underline{G(\mathbb{Q}_{p})}]$, by pulling back this HN-strata along $h^{\leftarrow}$. This defines a locally closed stratification of $[\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]$. Let $i_{b}:[\mathcal{F}\ell_{G,\mu^{-1}}^{b}/\underline{G(\mathbb{Q}_{p})}]\hookrightarrow[\mathcal{F}\ell_{G,\mu^{-1}}/\underline{G(\mathbb{Q}_{p})}]$ denote the associated locally closed immersion. We write $\pi_{\mathrm{HT}}^{b}:[\mathcal{S}^{b}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b}/\underline{G(\mathbb{Q}_{p})}]$ (resp. $\pi_{\mathrm{HT}}^{b,*}:[\mathcal{S}^{b,*}_{K_{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b}/\underline{G(\mathbb{Q}_{p})}]$) for the pullbacks of $\pi_{\mathrm{HT}}$ (resp. $\pi^{*}_{\mathrm{HT}}$) along $i_{b}$. On the good reduction locus, we also have an additional stratification coming from pulling back the Newton stratification on the special fiber along the specialization map. There is a rather subtle point that this does not agree with the pullback of the locally closed strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b}$ (namely, the closure relationships are opposite with respect to the partial ordering on $B(G)$). We write $\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}$ for these Newton strata coming from the special fiber. There exists a natural map
$\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}\times_{\mathcal{F}\ell_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b}\hookrightarrow\mathcal{S}^{b,\circ}_{K^{p},C}$
@@
which is a qcqs open immersion containing all rank $1$ points ([CS1, Page 68] [Ko, Page 8]). We write $\pi_{\mathrm{HT}}^{b,\circ}:[(\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}\times_{\mathcal{F}\ell_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b})/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b}/\underline{G(\mathbb{Q}_{p})}]$ for the Hodge-Tate period map on this locus, and similarly we write $\overline{\pi}_{\mathrm{HT}}^{b,\circ}:[\overline{\mathcal{S}}^{b,\circ}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b}/\underline{G(\mathbb{Q}_{p})}]$ for the induced map on the canonical compactification, where we note that this agrees with the canonical compactification of $[(\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}\times_{\mathcal{F}\ell_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b})/\underline{G(\mathbb{Q}_{p})}]$ by the previous remark on rank $1$ points and the fact that $\mathcal{F}\ell_{G,\mu^{-1}}^{b}$ is partially proper [^2] The partial properness of these strata follows directly from the moduli interpretation, since the category of vector bundles on the Fargues-Fontaine curve is insensitive to the ring of definition..
@@
Define the group diamond $\mathcal{J}_{b}:=\mathrm{Aut}(\mathcal{E}_{b})$, as in [FS, Proposition III.5.1]. We have an isomorphism $j_{b}:\mathrm{Bun}_{G}^{b}\simeq[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]\hookrightarrow\mathrm{Bun}_{G}$ with the locally closed HN-strata in $\mathrm{Bun}_{G}$ defined by $b$. There is a $\mathcal{J}_{b}$-torsor over the adic Newton strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b}$ given by rigidifying the bundle $\mathcal{E}$ to be isomorphic to $\mathcal{E}_{b}$. This gives a map
@@
$h^{\leftarrow}_{b}:[\mathcal{F}\ell_{G,\mu^{-1}}^{b}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]\simeq\mathrm{Bun}_{G,C}^{b}$

### v5

@@ 3.1. The Hodge-Tate Period Morphism
Observe that since both $\widehat{\mathscr{S}}(\mathbf{G},X)_{K^{p}}^{\mathrm{tor}}$ and the boundary $\widehat{\partial\mathscr{S}}(\mathbf{G},X)_{K^{p}}^{\mathrm{tor}}$ are proper, we have that the $\mathop{\rm Spa}(C^{\prime},O_{C^{\prime}})$ points of the adic generic fiber $\widehat{\mathscr{S}}(\mathbf{G},X)_{K^{p}}^{\mathrm{tor}}$ and $\mathcal{S}^{\mathrm{tor}}_{K^{p}}$ are the same, and similarly for the adic generic fiber of $\widehat{\partial\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K^{p}}$ and $\partial\mathcal{S}^{\mathrm{tor}}_{K^{p},C}$. Moreover, the $\mathop{\rm Spa}(C^{\prime},O_{C^{\prime}})$-points of $\widehat{\mathscr{S}}(\mathbf{G},X)_{K^{p}}^{\mathrm{tor}}$ are also the same as the $\mathrm{Spf}O_{C^{\prime}}$-points of the formal scheme $\widehat{\mathscr{S}}(\mathbf{G},X)_{K^{p},O_{C}}^{\mathrm{tor}}$ . A similar result also holds for the boundary. Thus, we can apply Proposition 2.13 to conclude, since $\tilde{\partial}\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}$ is defined as the fiber product of formal schemes $\mathfrak{Ig}^{b,\mathrm{tor}}_{O_{C}}\times_{\widehat{\mathscr{S}}(\mathbf{G},X)_{K^{p},O_{C}}}\widehat{\partial\mathscr{S}}(\mathbf{G},X)^{\mathrm{tor}}_{K^{p},O_{C}}$.
@@
For a geometric point $x:\mathop{\rm Spa}(C,O_{C})\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$, we have an identification:
@@
We have an identification $\pi_{\mathrm{HT}}^{-1}(x)\simeq\overline{\mathfrak{Ig}}^{b,*}_{C}\setminus\overline{\partial\mathfrak{Ig}}^{b,*}_{C}$ by the previous corollary, so, by proper base-change, we are tasked with computing the compactly supported cohomology of this space. We note, by Theorem 3.2, the adic spaces $\mathfrak{Ig}^{b,*}_{C}$ are affinoid perfectoid. It follows that the canonical compactification $\overline{\mathfrak{Ig}}^{b,*}_{C}\simeq(\pi_{\mathrm{HT}}^{*})^{-1}(x)$ is also affinoid perfectoid, by [56, Proposition 18.7 (iv)]. In particular, it is quasi-compact and partially proper, so in particular proper. It therefore follows by excision [^5] One easily checks that the excision sequence is exact on geometric points, and this is sufficient by [56, Proposition 14.3]. that we have a distinguished triangle
@@
Applying Theorem 3.1 again, we know that $k:\mathfrak{Ig}^{b,*}\hookrightarrow\overline{\mathfrak{Ig}}^{b,*}_{C}$ is a qcqs open immersion of perfectoid spaces inducing an isomorphism on rank $1$ points, and it follows that the same is true for the induced map on the Zariski closed subspaces $\phantom{}{}^{\partial}k:\partial\mathfrak{Ig}^{b,*}\hookrightarrow\overline{\partial\mathfrak{Ig}}^{b,*}_{C}$. Therefore, we can apply [7, Lemma 4.4.2], this tells us that the natural maps
@@
are isomorphisms, giving identifications $R\Gamma(\overline{\mathfrak{Ig}}_{C}^{b,*},\overline{\mathbb{F}}_{\ell})\simeq R\Gamma(\mathfrak{Ig}^{b,*}_{C},\overline{\mathbb{F}}_{\ell})$ and $R\Gamma(\overline{\partial\mathfrak{Ig}}_{C}^{b,*},\overline{\mathbb{F}}_{\ell})\simeq R\Gamma(\partial\mathfrak{Ig}^{b,*}_{C},\overline{\mathbb{F}}_{\ell})$. Now, by [7, Lemma 4.4.3], we have further identifications of $R\Gamma(\partial\mathfrak{Ig}^{b,*}_{C},\overline{\mathbb{F}}_{\ell})$ and $R\Gamma(\mathfrak{Ig}^{b,*}_{C},\overline{\mathbb{F}}_{\ell})$ with the cohomology of the perfect schemes $\partial\mathrm{Ig}^{b,*}$ and $\mathrm{Ig}^{b,*}$, respectively. Substituting this into the triangle (6), we get a distinguished triangle
@@
By applying quasi-compact base-change [56, Proposition 17.6] and then using that the inclusion $\partial\mathfrak{Ig}_{C}^{b,*}\subset\mathfrak{Ig}_{C}^{b,*}$ is induced from taking the rigid generic fiber over $C$ of Witt vectors applied to $\partial\mathrm{Ig}^{b,*}\subset\mathrm{Ig}^{b,*}$, we identify the last map with the natural restriction map on the cohomology. However, this identifies the first term with precisely the partially compactly supported cohomology, as desired. ∎
@@
Let $d_{b}:=\langle 2\rho_{G},\nu_{b}\rangle$. We have $d_{b}=\mathrm{dim}(\mathrm{Ig}^{b,*})=\mathrm{dim}(\mathrm{Ig}^{b})$, and the cohomology of the complex
@@
From [23, Corollary 7.8], we know that the dimension of $\mathscr{C}_{\mathbb{X}_{b}}$ is $d_{b}$, and since
$\mathrm{Ig}^{b}\rightarrow\mathscr{C}_{\mathbb{X}_{b}}^{\mathrm{perf}},$
is a pro-étale cover with Galois group $\mathrm{Aut}_{G}(\mathbb{X})(\overline{\mathbb{F}}_{p})$ over the perfection of the central leaf attached to $\mathbb{X}_{b}$, we may conclude the first part about dimensions. We also saw in Theorem 3.2 that $\mathrm{Ig}^{b,*}$ is an affine scheme. So we would like to apply Artin vanishing [1, Exposé XIV, Corrolary 3.2]; however, $\mathrm{Ig}^{b,*}$ is also a perfect scheme so not in general of finite type. Recall from [7, Proposition 4.3.8] that $\mathrm{Ig}^{b}$ is obtained as the perfection of the limit of the finite étale covers
$\mathrm{Ig}^{b}_{m}\rightarrow\mathscr{C}_{\mathbb{X}_{b}},$
where $\mathrm{Ig}^{b}_{m}$ is as defined in [8, Remark 2.3.6]. These spaces are of finite type over $\overline{\mathbb{F}}_{p}$. Define $\mathrm{Ig}^{b,*}_{m}$ to be the normalization of $\mathscr{C}_{\mathbb{X}}^{*}$ in $\mathrm{Ig}^{b}_{m}$ of the finite étale cover $\mathrm{Ig}^{b}_{m}\rightarrow\mathscr{C}_{\mathbb{X}_{b}}$. By [54, Theorem 3.33], $\mathscr{C}_{\mathbb{X}}^{*}$ is affine of dimension $d_{b}$; therefore, it follows that $\mathrm{Ig}^{b,*}_{m}$ is a normal and affine scheme of dimension $d_{b}$ which will be of finite type, since $\mathrm{Ig}^{b}_{m}$ is. Let $g_{b,m}:\mathrm{Ig}^{b}_{m}\rightarrow\mathrm{Ig}^{b,*}_{m}$ be the natural map, then Artin vanishing implies that the cohomology of the complex $g_{b,m!}(\overline{\mathbb{F}}_{\ell})$ is concentrated in degrees $\leq d_{b}$. By definition of $\mathrm{Ig}^{b,*}$, it is the perfection of $\mathop{\rm lim}\nolimits_{m\geq 1}\mathrm{Ig}_{m}^{b,*}$. Therefore, since étale cohomology is invariant under perfection, we can conclude by applying [61, Tag 09QY] to the system of sheaves $g_{b,m!}(\overline{\mathbb{F}}_{\ell})$. ∎
Now we would like to link this analysis with the semi-perversity of certain sheaves on $\mathrm{Bun}_{G}$ via a version of the Mantovan product formula.
3.2. Mantovan’s Product Formula
We consider the Hodge-Tate period morphism
$\pi_{\mathrm{HT}}:[\mathcal{S}_{K^{p}}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$
from the previous section quotiented out by $G(\mathbb{Q}_{p})$, where we abusively use $\pi_{\mathrm{HT}}$ for this quotiented out map. We let $h^{\rightarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\underline{G(\mathbb{Q}_{p})}]\simeq\mathrm{Bun}_{G,C}^{1}$ be the structure map quotiented out by $G(\mathbb{Q}_{p})$. Note this is a proper map, since $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$ is proper over $\mathop{\rm Spd}(C)$.
@@ 3.2. Mantovan’s Product Formula
(7) $R\Gamma_{c}(\mathcal{S}_{K^{p},C},\overline{\mathbb{F}}_{\ell})\simeq Rh^{\rightarrow}_{*}R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$
of $G(\mathbb{Q}_{p})$-representations.
Remark 3.7.
@@
$h^{\leftarrow}:[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow\mathrm{Bun}_{G}$
@@
remembering the isomorphism class of the bundle $\mathcal{E}$ in the moduli interpretation of $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$ as a diamond described in §3.1. This defines a cohomologically smooth map by [16, Theorem IV.1.19], and the image identifies with the open subset $B(G,\mu)\subset B(G)$ under the identification $|\mathrm{Bun}_{G}|\simeq B(G)$ of topological spaces [65], where $|\mathrm{Bun}_{G}|$ denotes the underlying topological space of the $v$-stack $\mathrm{Bun}_{G}$ and $B(G)$ has the topology given by its natural partial ordering (i.e it is generated by the opens $U_{b}:=\{x\in B(G)|x\leq b\}$, where $x\leq b$ means that their Kottwitz invariants $\kappa(x)=\kappa(b)$ agree and the differences of slopes $\nu_{b}-\nu_{x}$ is given by a $\mathbb{Q}_{\geq 0}$-linear combination of positive coroots (See [16, Section I.5] for details)). For each $b$, we have a locally closed Harder-Narasimhan stratum $j_{b}:\mathrm{Bun}_{G,C}^{b}\hookrightarrow\mathrm{Bun}_{G,C}$, and we can define the locally closed subset $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$, by pulling back this HN-strata along $h^{\leftarrow}$. This defines a locally closed stratification of $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$. Let $i_{b}:[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]\hookrightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ denote the associated locally closed immersion. We write $\pi_{\mathrm{HT}}^{b}:[\mathcal{S}^{b}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ (resp. $\pi_{\mathrm{HT}}^{b,*}:[\mathcal{S}^{b,*}_{K_{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$) for the pullbacks of $\pi_{\mathrm{HT}}$ (resp. $\pi^{*}_{\mathrm{HT}}$) along $i_{b}$. On the good reduction locus, we also have an additional stratification coming from pulling back the Newton stratification on the special fiber along the specialization map. We recall the rather subtle point that this does not agree with the pullback of the locally closed strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$ (namely, the closure relationships are opposite with respect to the partial ordering on $B(G)$). We write $\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}$ for these Newton strata coming from the special fiber. There exists a natural map
$\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}\times_{\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}}\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}\hookrightarrow\mathcal{S}^{b,\circ}_{K^{p},C}$
@@
which is a qcqs open immersion containing all rank $1$ points ([7, Page 68], [37, Page 8]). We write $\pi_{\mathrm{HT}}^{b,\circ}:[(\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond})/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ for the Hodge-Tate period map on this locus, and similarly we write $\overline{\pi}_{\mathrm{HT}}^{b,\circ}:[\overline{\mathcal{S}}^{b,\circ}_{K^{p},C}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]$ for the induced map on the canonical compactification, where we note that this agrees with the canonical compactification of $(\mathcal{S}^{b,\circ,\mathrm{rd}}_{K^{p},C}\times_{\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}}\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond})/\underline{G(\mathbb{Q}_{p})}]$ by the previous remark on rank $1$ points and the fact that $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$ is partially proper [^6] The partial properness of these strata follows directly from the moduli interpretation, since the category of vector bundles on the Fargues-Fontaine curve is insensitive to the choice of $R^{+}$ for a Tate Huber pair $(R,R^{+})$..
@@
Define the group diamond $\mathcal{J}_{b}:=\mathrm{Aut}(\mathcal{E}_{b})$, as in [16, Proposition III.5.1] parametrizing automorphism of the $G$-bundle $\mathcal{E}_{b}$ on the Fargues-Fontaine curve attached to $b$. We have an isomorphism $j_{b}:\mathrm{Bun}_{G}^{b}\simeq[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]\hookrightarrow\mathrm{Bun}_{G}$ with the locally closed HN-strata in $\mathrm{Bun}_{G}$ defined by $b$. There is a $\mathcal{J}_{b}$-torsor over the adic Newton strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$ given by rigidifying the bundle $\mathcal{E}$ to be isomorphic to $\mathcal{E}_{b}$. This gives a map
@@
(8) $h^{\leftarrow}_{b}:[\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}/\underline{G(\mathbb{Q}_{p})}]\rightarrow[\mathop{\rm Spd}(C)/\mathcal{J}_{b}]\simeq\mathrm{Bun}_{G,C}^{b}$